# Foundational Policy Acquisition via Multitask Learning for Motor Skill Generation

### v3

@@ Acknowledgments
This study used several opensource software projects, such as PyTorch [30], rlkit [31], dm-control [32], and Optuna [27]. We are grateful to the project contributors for providing valuable machine learning tools. This work was supported by JSPS KAKENHI (Grant number: 19J22987); by project JPNP20006, commissioned by NEDO; by JST Moonshot R&D program (Grant number: JPMJMS223B-3); and by Tateisi Science and Technology Foundation.
@@
We adopted the target networks, clipped double-$Q$ learning, and entropy autotuning in the policy generation phase. The target network and clipped double-$Q$ learning are the methods for reducing the estimation error of $Q$ functions. For this purpose, four neural networks were trained to approximate a single $Q$ function. Entropy autotuning changes the temperature parameter $\alpha$ that controls the exploration and exploitation trade-off.
@@
$Q$-function networks were trained to reduce Bellman residuals, as shown in Eq. (17) as follows:
@@
where $\phi_{i}$ and $\phi_{\mathrm{target}}$ are the $Q$-network parameters; $r$, $\gamma$, and $\alpha$ denote the reward function, discount factor, and temperature parameter, respectively; $\pi$ represents the policy; $q$ represents the encoder; and $\mathcal{D}$ is the state-action pair data stored in a replay buffer. The value function $\bar{V}$ was computed from $Q(s^{\prime},a^{\prime},c)$, where $s^{\prime}$, $a^{\prime}$ and $c$ are the state, action, and context at the next step, respectively. Simply using the above method tends to overestimate the action values; therefor, clipped double-$Q$ Learning was proposed to overcome this overestimation problem [33]. To use this method, we trained two networks and selected the smaller output of the two approximated $Q$ functions to derive the value $\bar{V}$. In practice, $Q$ values at the next step are derived from the target network $\phi_{\mathrm{target}}$ explained below.
@@
The target networks were used to stabilize the learning process of the $Q$ function. As we explained in the previous section, the target value $V$ was derived from the $Q$ network, which caused the target value estimation to be unstable [34]. Therefore, as suggested in [35], we updated the target network through moving average as follows:
@@
To autotune the temperature parameter $\alpha$ of the entropy regularization term, we considered the following optimization problem [36]: $H$ is the conditional entropy of the policy distribution given the state $s$, and $\bar{H}$ is the target entropy set by an experimenter. The target entropy was set to $-1$ times the action dimension $|A|$ as follow [36]:
@@
A heuristic transformation $\alpha\leftarrow\exp(\alpha)$ was performed on the temperature parameter $\alpha$ to guarantee the positive value as practical implementation  [^1] https://github.com/rail-berkeley/softlearning/blob/master/softlearning/algorithms/sac.py#L140. Accessed: July-10, 2023..
@@
The parameters used in this study are listed in Table III. Notably, most of the parameters were selected based on those reported in the study by Haarnoja et al.[36].
@@
hyperparameter policy search adaptation
action time period 25ms 25 ms
neural network optimizer Adam Adam
Policy and encoder learning rate $3\times 10^{-4}$
@@
collect step size per training 1000 100
collect step size per evaluation 1000 100
@@
optimization algorithm Adam Adam
@@
$k_{\mathrm{max}}$ 100

### v4

@@ Acknowledgments
This work was supported by JSPS KAKENHI (Grant number: JP19J22987, JP22H04998, JP23K24925); by project JPNP20006, commissioned by NEDO; by JST Moonshot R&D program (Grant number: JPMJMS223B-3); and by Tateisi Science and Technology Foundation.
@@
We adopted the target networks, clipped double-$Q$ learning, and entropy autotuning in the foundational policy acquisition.
@@
double $Q$-function networks were trained to reduce Bellman residuals, as shown in Eq. (19) as follows:
@@
where $\phi_{i}$ and $\phi_{\mathrm{target}}$ are the $Q$-network parameters; $r$, $\gamma$, and $\alpha$ denote the reward function, discount factor, and temperature parameter, respectively; $\pi$ represents the policy; $q$ represents the encoder; and $\mathcal{D}$ is the state-action pair data stored in a replay buffer. The value function $\bar{V}$ was computed from $Q(s^{\prime},a^{\prime},c)$, where $s^{\prime}$, $a^{\prime}$ and $c$ are the state, action, and context at the next step, respectively. Simply using the above method tends to overestimate the action values; therefor, clipped double-$Q$ Learning was proposed to overcome this overestimation problem [39]. To use this method, we trained two networks and selected the smaller output of the two approximated $Q$ functions to derive the value $\bar{V}$.
@@
The target networks $\phi_{\mathrm{target}}$ were used to stabilize the learning process of the $Q$ function. As we explained in the previous section, the target value $V$ was derived from the $Q$ network, which caused the target value estimation to be unstable [40]. Therefore, as suggested in [41], we updated the target network through moving average as follows:
@@
To autotune the temperature parameter $\alpha$ of the entropy regularization term, we considered the following optimization problem [42]: $H$ is the conditional entropy of the policy distribution given the state $s$, and $\bar{H}$ is the target entropy set by an experimenter.
@@
The temperature parameter $\alpha$ is transformed $\alpha\leftarrow\exp(\alpha)$ to guarantee the positive value as practical implementation  [^1] https://github.com/rail-berkeley/softlearning/blob/master/softlearning/
algorithms/sac.py#L140. Accessed: July-10, 2023.
@@
The parameters used in this study are listed in Table III. Notably, most parameters are based on the original SAC [42].
@@
hyperparameter policy search adaptation info
action time period 50 / 25 ms 50 / 25 ms Cheetah / Monopod
@@
collect step size per training 1000 200/100 Cheetah / Monopod
collect step size per evaluation 200/1000 200/100 Cheetah / Monopod
@@
optimization algorithm Adam TPE
@@
target entropy $\bar{H}$ $-|A|$
number of candidates 5 / 25 - Cheetah / Monopod
skill generation iterations $k_{\mathrm{max}}$ 100